OpenStudy (anonymous):

Which shows three ordered pairs that are all solutions of the equation y = x + 7? A. (0, 7), (2, 9), and (3, 11) B. (0, 7), (2, 9), and (3, 10) C. (1, 7), (2, 9), and (3, 10) D. (1, 8), (2, 9), and (3, 11)

OpenStudy (anonymous):

Please Help.

OpenStudy (decarr432):

okay do you know how this is graphed?

OpenStudy (campbell_st):

just use substitution. You are being asked to identify the set of 2 points that are on the line y = x + 7 each ordered pair is in the form (x, y) so substitute each x value to see if you get the y= value e.g (0, 7) substitute x = 0 to see if you get y = 7 y = 0 + 7 so y = 7 so now you know the point (0,7) is on the line/ hope it helps